OpenStudy (anonymous):

The total cost of Mel’s cell phone was $230 after 5 mos of service and $390 after 9 mos. What costs had Mel already incurred when his service just began?

OpenStudy (anonymous):

(390-230)/(9-5) = cost per month 160/4 = $40/month 230 - 5* 40 = 30 dollars initial costs